### v1

# The convexity of a planar domain via properties of solutions to the modified Helmholtz equation

The convexity of a planar domain via properties of solutions
to the modified Helmholtz equation
Nikolay Kuznetsov
Laboratory for Mathematical Modelling of Wave Phenomena Institute for Problems in Mechanical Engineering, RAS V.O., Bol’shoy pr. 61, St Petersburg 199178, Russia nikolay.g.kuznetsov@gmail.com
A new characterization of convexity of a planar domain is obtained. Its derivation involves two classical facts: the Varadhan’s formula, expressing the distance function with respect to the domain’s boundary via real-valued solutions of the modified Helmholtz equation, and the convexity of a planar domain in which the distance function is superharmonic.
1 Introduction
Let a point $x=(x_{1},x_{2})\in\mathbb{R}^{2}$ belongs to a half-plane, say $x_{2}>0$. It is obvious that the distance from $x$ to the $x_{1}$-axis bounding this half-plane — the length of the segment orthogonal to the $x_{1}$-axis is equal to $x_{2}$ — is a harmonic function in this domain. The converse assertion is far from being trivial to prove; it is referred to as one of “three secrets about harmonic functions” in the interesting note [4]. For an arbitrary planar domain $D$ the distance function with respect to $\partial D$ is
$d(x,\partial D)=\inf_{y\in\partial D}|x-y|\,,$
where $x$ is a point in $D$ and $|x-y|=\sqrt{(x_{1}-y_{1})^{2}+(x_{2}-y_{2})^{2}}$ is the Euclidean norm in $\mathbb{R}^{2}$. The superharmonicity of this function in $D$ implies that this domain is convex; see Theorem 1.3 below (it was obtained by Armitage and Kuran [3] in 1985). In the present note, it is shown that real-valued solutions of the modified Helmholtz equation (they are labelled panharmonic functions in [5]) also provide a characterization of convexity.
1.1. Statement of the problem; formulation of the result. Throughout this paper $D$ denotes a bounded domain of the Euclidean plane $\mathbb{R}^{2}$. Its boundary $\partial D$ is assumed to be sufficiently smooth, say $C^{2}$, and so the Dirichlet problem for the modified Helmholtz equation
$\nabla^{2}v-\mu^{2}v=0\ \ \mbox{for}\ x=(x_{1},x_{2})\in D,\quad\mu\in\mathbb{R}_{+},$ (1.1)
has a unique solution satisfying the boundary condition
$v(x)=1\quad\mbox{for}\ x\in\partial D;$ (1.2)
here and below $\nabla=(\partial_{1},\partial_{2})$ is the gradient operator, $\partial_{i}=\partial/\partial x_{i}$. It is known (see [8, p. 437]) that a solution of this problem, say $v(x,\mu)$, satisfies the inequalities
$00\leqslant v(x,\mu)\leqslant 1\ \ \mbox{for all}\ x\in\overline{D}\ \mbox{and any}\ \mu>$ 0 . (1.3)
It occurs that problem’s solutions corresponding to large values of $\mu$ characterize the domain $D$. Namely, our aim is to prove the following
Theorem 1.1.
Let $D\subset\mathbb{R}^{2}$ be a bounded domain with smooth boundary. If the conditions
$v(x,\mu)\neq 0\ \ and\ \ |\nabla v(x,\mu)|\leqslant\mu v(x,\mu),\ \ x\in D,$ (1.4)
are fulfilled for solutions of problem (1.1), (1.2) with all large $\mu>0$, then $D$ is convex.
This theorem is proved in Section 2, and here we illustrate it with an elementary example in which the domain is a disc, and so convex.
Let $I_{0}$ be the modified Bessel function of the first kind (see, for example, [5, p. 111]), then $v(x,\mu)=aI_{0}(\mu|x|)$ with $a\in(0,1)$ solves problem (1.1), (1.2) in the disc $B_{r}(0)=\{y\in\mathbb{R}^{2}:|y|<r\}$ provided $r>0$ is such that $aI_{0}(\mu r)=1$.
It is clear that properties of the function $I_{0}$ (see [2, sect. 9.6]) yield the first condition (1.4) for $v(\cdot,\mu)$, whereas the second condition follows from the inequality $I_{1}(\mu|x|)<I_{0}(\mu|x|)$, because
$|\nabla v(x,\mu)|=\mu aI_{1}(\mu|x|)$
in view of formula [2, 9.6.27].
1.2. Background. Here we formulate two classical results about the distance function with respect to $\partial D$. These results are essential for proving Theorem 1.1.
Theorem 1.2 (Varadhan, [8]).
Let $D\subset\mathbb{R}^{2}$ be a bounded domain with smooth boundary. If $v(x,\mu)$ is a solution of the Dirichlet problem (1.1), (1.2), then
$-\mu^{-1}\log v(x,\mu)\to d(x,\partial D)\quad as\ \mu\to\infty$ (1.5)
uniformly on $D$.
It is worth mentioning that the original Varadhan’s result is valid for more general elliptic equations, but its proof is rather complicated. In the recent preprint [6], Theorem 1.2 is generalized by adding the source term to equation (1.1); moreover, the rate of convergence of $-\mu^{-1}\log v(x,\mu)$ to $d(x,\partial D)$ is evaluated. The next result has a simple proof which is reproduced here.
Theorem 1.3 (Armitage and Kuran, [3]).
If $d(\cdot,\partial D)$ is superharmonic in a domain $D\subset\mathbb{R}^{2}$, then $D$ is convex.
Proof.
Let us assume $D\subset\mathbb{R}^{2}$ to be nonconvex and show that $d(\cdot,\partial D)$ is not superharmonic in $D$. According to the assumption, there exist a point $y\in\partial D$, a closed half-plane $P$ with $y\in\partial P$ (without loss of generality, we set $y=(0,0)$ and $P=\{x\in\mathbb{R}^{2}:x_{2}\geqslant 0\}$), and $r>0$ such that
$P\cap(\overline{B_{r}(y)}\setminus\{y\})\subset D\,,\ \ \mbox{whereas}\ \ (\partial D\setminus\{y\})\cap B_{r}(y)\subset\mathbb{R}^{2}\setminus P\,,$
where $B_{r}(y)=\{x\in\mathbb{R}^{2}:|x-y|<r\}$ (cf. [7, Theorem 4.8]).
Then for $x\in B=B_{r/8}(x_{0})$, where $x_{0}=(0,r/4)$, we have that $d(x,\partial D)>x_{2}$ provided $x_{1}\neq 0$. Hence
[…]
and so the area mean-value inequality, guaranteeing the superharmonicity of $d(\cdot,\partial D)$, is violated for $B$. ∎
Theorem 1.3 is specifically a two-dimensional one because neither $D$ nor $\overline{D}$ need be convex in higher dimensions; an example confirming this is given in [3, Sect. 5].
2 Proof of Theorem 1.1
Let us show that conditions (1.4) imply the superharmonicity of $d(\cdot,\partial D)$ in $D$, which allows us to apply Theorem 1.3, thus yielding the convexity of $D$.
For a nonvanishing $u\in C^{2}(D)$, one obtains by a straightforward calculation that
$\nabla^{2}(\log u)=-\frac{|\nabla u|^{2}}{u^{2}}+\frac{\nabla^{2}u}{u}\quad\mbox{in}\ D.$
Therefore, if $v(\cdot,\mu)$ is a solution of the Dirichlet problem (1.1), (1.2) with large $\mu>0$, then
$-\nabla^{2}(\log v)=-\mu^{2}+|\nabla v|^{2}/v^{2}<0$
provided conditions (1.4) are valid. Hence, $-\mu^{-1}\log v(\cdot,\mu)$ is superharmonic in $D$ for all large $\mu>0$.
The next step is to demonstrate that passage to the limit in Theorem 1.2 preserves superharmonicity. The proof is based on the following corollary of the Varadhan’s result [8, Theorem 3.6].
Proposition 2.1.
Let $v(x,\mu)$ be a solution of the Dirichlet problem (1.1), (1.2) in a bounded domain $D\subset\mathbb{R}^{2}$ with smooth boundary. Then for every $\rho\in(0,1/2)$, there exists a constant $C_{\rho}>1$ such that the estimate
$v(x,\mu)\leqslant C_{\rho}\exp\{-\mu\,(1-\rho)\,d(x,\partial D)\}\,,\quad x\in\overline{D},$
is valid for all $\mu>0$.
In view of inequality (1.3) and the first condition (1.4), this assertion implies that
$-\mu^{-1}\log v(x,\mu)\geqslant-\mu^{-1}\log C_{\rho}+(1-\rho)\,d(x,\partial D)\,,\quad x\in\overline{D}.$ (2.1)
Averaging this over an arbitrary open disc $B_{r}(x)$ such that $\overline{B_{r}(x)}\subset D$, we obtain
$-\mu^{-1}\log v(x,\mu)\geqslant-\mu^{-1}\log C_{\rho}+\frac{1-\rho}{\pi r^{2}}\int_{B_{r}(x)}d(y,\partial D)\,\mathrm{d}\kern 0.2pty\,,$
because the function on the left-hand side of (2.1) is superharmonic and the first term on the right is harmonic. Letting $\mu\to\infty$ first and then $\rho\to+0$, we arrive at
$d(x,\partial D)\geqslant\frac{1}{\pi r^{2}}\int_{B_{r}(x)}d(y,\partial D)\,\mathrm{d}\kern 0.2pty\,,\quad x\in D,$
in view of (1.5).
Since the obtained inequality is valid for all $r>0$ such that $\overline{B_{r}(x)}\subset D$, the distance function $d(\cdot,\partial D)$ is superharmonic in $D$. Then Theorem 1.3 yields the assertion of Theorem 1.1, and this completes the proof.