### v2

# The convexity of a planar domain via properties of solutions to the modified Helmholtz equation

The convexity of a planar domain via properties of solutions
to the modified Helmholtz equation
Nikolay Kuznetsov
Laboratory for Mathematical Modelling of Wave Phenomena Institute for Problems in Mechanical Engineering, RAS V.O., Bol’shoy pr. 61, St Petersburg 199178, Russia nikolay.g.kuznetsov@gmail.com
A new characterization of convexity of a planar domain is obtained. Its derivation involves two classical facts: the Varadhan’s formula, expressing the distance function with respect to the domain’s boundary via real-valued solutions of the modified Helmholtz equation, and the convexity of a planar domain in which the distance function is superharmonic.
1 Introduction
Let a point $x=(x_{1},x_{2})\in\mathbb{R}^{2}$ belong to a half-plane, say $x_{2}>0$. It is obvious that the distance from $x$ to the $x_{1}$-axis bounding this half-plane — the length of the segment orthogonal to the $x_{1}$-axis — is equal to $x_{2}$, and so is harmonic in this domain. The converse assertion is far from being trivial to prove; it is referred to as one of “three secrets about harmonic functions” in the interesting note [4]. A half-plane is the maximal, as it were, convex domain in $\mathbb{R}^{2}$, whereas our aim is to characterize in a new fashion the convexity of a bounded planar domain.
In an arbitrary planar domain $D$, the distance function with respect to $\partial D$ is
$d(x,\partial D)=\inf_{y\in\partial D}|x-y|\,,$
where $x\in D$ and $|x-y|=\sqrt{(x_{1}-y_{1})^{2}+(x_{2}-y_{2})^{2}}$ is the Euclidean norm in $\mathbb{R}^{2}$. It is known that the superharmonicity of this function in $D$ implies that the domain is convex; see Theorem 1.3 below (it was obtained by Armitage and Kuran [3] in 1985). In the present note, we demonstrate that it is also possible to describe the domain’s convexity by means of real-valued solutions of the modified Helmholtz equation (labelled panharmonic functions in [5]); indeed, this is achieved by imposing some conditions on them.
1.1. Statement of the problem; formulation of the result. Throughout this paper $D$ denotes a bounded domain of the Euclidean plane $\mathbb{R}^{2}$. Its boundary $\partial D$ is assumed to be sufficiently smooth, say $C^{2}$, and so the Dirichlet problem for the modified Helmholtz equation
$\nabla^{2}v-\mu^{2}v=0\ \ \mbox{for}\ x=(x_{1},x_{2})\in D,\quad\mu\in\mathbb{R}_{+},$ (1.1)
has a unique solution satisfying the boundary condition
$v(x)=1\quad\mbox{for}\ x\in\partial D;$ (1.2)
here and below $\nabla=(\partial_{1},\partial_{2})$ is the gradient operator, $\partial_{i}=\partial/\partial x_{i}$.
It has long bean known (see [8, p. 437]) that a solution of this problem, say $v(x,\mu)$, satisfies
$00\leqslant v(x,\mu)\leqslant 1\ \ \mbox{for all}\ x\in\overline{D}\ \mbox{and any}\ \mu>$ 0 . (1.3)
Moreover, the left inequality is strict; see [6, Sect. 2.2] for the proof.
It occurs that problem’s solutions corresponding to large values of $\mu$ characterize the domain $D$. Namely, our aim is to prove the following
Theorem 1.1.
Let $D\subset\mathbb{R}^{2}$ be a bounded domain with smooth boundary. If the condition
$|\nabla v(x,\mu)|\leqslant\mu v(x,\mu),\ \ x\in D,$ (1.4)
holds for solutions of problem (1.1), (1.2) with all large $\mu>0$, then $D$ is convex.
This theorem is proved in Section 2, and here we illustrate it with two elementary examples.
In the first one, the domain $D$ is the half-plane $x_{2}>0$, in which case it is natural to require
$v(x,\mu)\to 0\quad\mbox{as}\ x_{2}\to\infty.$
Duffin established that $v(x,\mu)=\mathrm{e}\kern 0.2pt^{-\mu x_{2}}$ solves problem (1.1), (1.2) complemented by the latter condition; see his proof of [5, Theorem 5]. Condition (1.4) is fulfilled for this function for all $\mu>0$, and so Theorem 1.1 implies the obvious fact that $D$ is convex. Notice that
$-\mu^{-1}\log v(x,\mu)=x_{2}=d(x,\partial D)$
in agreement with the assertion of Theorem 1.2 formulated below.
In the second example, the domain $D$ is a disc, and so convex. As usual, $I_{0}$ stands for the modified Bessel function of the first kind (see, for example, [5, p. 111]). If $a\in(0,1)$, then $v(x,\mu)=aI_{0}(\mu|x|)$ solves problem (1.1), (1.2) in the disc $B_{r}(0)=\{y\in\mathbb{R}^{2}:|y|<r\}$ provided $r>0$ is such that $aI_{0}(\mu r)=1$.
It is clear that properties of the function $I_{0}$ (see [2, sect. 9.6]) yield condition (1.4) for $v(\cdot,\mu)$; indeed, it takes the form $I_{1}(\mu|x|)<I_{0}(\mu|x|)$, because
$|\nabla v(x,\mu)|=\mu aI_{1}(\mu|x|)$
in view of formula [2, 9.6.27].
1.2. Background. Here we formulate two classical results about the distance function with respect to $\partial D$. These results are essential for proving Theorem 1.1.
Theorem 1.2 (Varadhan, [8]).
Let $D\subset\mathbb{R}^{2}$ be a bounded domain with smooth boundary. If $v(x,\mu)$ is a solution of the Dirichlet problem (1.1), (1.2), then
$-\mu^{-1}\log v(x,\mu)\to d(x,\partial D)\quad as\ \mu\to\infty$ (1.5)
uniformly on $D$.
It is worth mentioning that the original Varadhan’s result involves the equation in $\mathbb{R}^{m}$, $m\geqslant 2$, with a general elliptic operator of second order instead of the Laplacian, and the distance is induced by a Riemannian metric derived from the operator’s coefficients. For this reason the Varadhan’s proof is rather complicated. In the recent preprint [6], signed distance functions are considered, thus generalizing $d(\cdot,\partial D)$. For this purpose the source term is added to equation (1.1), which allows the authors to extend Theorem 1.2; moreover, the rate of convergence of $-\mu^{-1}\log v(x,\mu)$ to $d(x,\partial D)$ is evaluated in the preprint.
Furthermore, let us consider $v(x,\mu)$, satisfying equation (1.1) and the Neumann condition
$\partial v/\partial n=\mu\quad\mbox{on}\ \partial D$
($n$ is the exterior normal on $\partial D$), instead of the Dirichlet condition (1.2). It is interesting to find out whether $-\mu^{-1}\log v(x,\mu)$ converges to $d(x,\partial D)$ in $D$ as $\mu\to\infty$ in this case.
Another classical result has a simple proof which is reproduced here.
Theorem 1.3 (Armitage and Kuran, [3]).
If $d(\cdot,\partial D)$ is superharmonic in a domain $D\subset\mathbb{R}^{2}$, then $D$ is convex.
Proof.
Let us assume $D\subset\mathbb{R}^{2}$ to be nonconvex and show that $d(\cdot,\partial D)$ is not superharmonic in $D$. According to the assumption, there exist a point $y\in\partial D$, a closed half-plane $P$ with $y\in\partial P$ (without loss of generality, we set $y=(0,0)$ and $P=\{x\in\mathbb{R}^{2}:x_{2}\geqslant 0\}$), and $r>0$ such that
$P\cap(\overline{B_{r}(y)}\setminus\{y\})\subset D\,,\ \ \mbox{whereas}\ \ (\partial D\setminus\{y\})\cap B_{r}(y)\subset\mathbb{R}^{2}\setminus P\,,$
where $B_{r}(y)=\{x\in\mathbb{R}^{2}:|x-y|<r\}$ (cf. [7, Theorem 4.8]).
Then for $x\in B=B_{r/8}(x_{0})$, where $x_{0}=(0,r/4)$, we have that $d(x,\partial D)>x_{2}$ provided $x_{1}\neq 0$. Hence
[…]
and so the area mean-value inequality, guaranteeing the superharmonicity of $d(\cdot,\partial D)$, is violated for $B$. ∎
It should be emphasized that Theorem 1.3 is specifically a two-dimensional result, because neither $D$ nor $\overline{D}$ need be convex in higher dimensions; in [3, Sect. 5], Armitage and Kuran give an example confirming the latter assertion.
2 Proof of Theorem 1.1
It is sufficient to show that conditions (1.4) imply that $d(\cdot,\partial D)$ is superharmonic in $D$; indeed, this allows us to apply Theorem 1.3, thus demonstrating the convexity of $D$.
For a nonvanishing $u\in C^{2}(D)$, one obtains by a straightforward calculation that
$\nabla^{2}(\log u)=-\frac{|\nabla u|^{2}}{u^{2}}+\frac{\nabla^{2}u}{u}\quad\mbox{in}\ D.$
Therefore, if $v(\cdot,\mu)$ is a solution of the Dirichlet problem (1.1), (1.2) with large $\mu>0$, then it does not vanish, and so
$-\nabla^{2}(\log v)=-\mu^{2}+|\nabla v|^{2}/v^{2}<0$
provided condition (1.4) is valid. Hence, $-\mu^{-1}\log v(\cdot,\mu)$ is superharmonic in $D$ for all large $\mu>0$.
The next step is to demonstrate that passage to the limit in Theorem 1.2 preserves superharmonicity. The proof is based on the following corollary of the Varadhan’s result [8, Theorem 3.6].
Proposition 2.1.
Let $v(x,\mu)$ be a solution of the Dirichlet problem (1.1), (1.2) in a bounded domain $D\subset\mathbb{R}^{2}$ with smooth boundary. Then for every $\rho\in(0,1/2)$, there exists a constant $C_{\rho}>1$ such that the estimate
$v(x,\mu)\leqslant C_{\rho}\exp\{-\mu\,(1-\rho)\,d(x,\partial D)\}\,,\quad x\in\overline{D},$
is valid for all $\mu>0$.
In view of inequality (1.3) and the first condition (1.4), this assertion implies that
$-\mu^{-1}\log v(x,\mu)\geqslant-\mu^{-1}\log C_{\rho}+(1-\rho)\,d(x,\partial D)\,,\quad x\in\overline{D}.$ (2.1)
Averaging this over an arbitrary open disc $B_{r}(x)$ such that $\overline{B_{r}(x)}\subset D$, we obtain
$-\mu^{-1}\log v(x,\mu)\geqslant-\mu^{-1}\log C_{\rho}+\frac{1-\rho}{\pi r^{2}}\int_{B_{r}(x)}d(y,\partial D)\,\mathrm{d}\kern 0.2pty\,,$
Indeed, the function on the left-hand side of (2.1) is superharmonic and the first term on the right is harmonic. Letting $\mu\to\infty$ first and then $\rho\to+0$, we see that the last inequality turns into
$d(x,\partial D)\geqslant\frac{1}{\pi r^{2}}\int_{B_{r}(x)}d(y,\partial D)\,\mathrm{d}\kern 0.2pty\,,\quad x\in D,$
in view of (1.5).
Since the obtained inequality is valid for all $r>0$ such that $\overline{B_{r}(x)}\subset D$, the distance function $d(\cdot,\partial D)$ is superharmonic in $D$. Then Theorem 1.3 yields the assertion of Theorem 1.1, thus completing the proof.